OpenStudy (anonymous):

PLEASE HELP!!!! If ABC~XYZ, what is the measure of a? Let c=12, z=18 and x=27 😩

OpenStudy (campbell_st):

well if the triangles are similar then the ratios of the corresponding sides are the same so \[\frac{a}{x} = \frac{c}{z}\] using the values \[\frac{a}{27} = \frac{12}{18}\] just solve for x

OpenStudy (anonymous):

Ohh ok thank you! ;))